OpenStudy (anonymous):

A country's population in 1995 was 228 million. In 2001 it was 230 million. Estimate the population in 2010 using the exponential growth formula. Round your answer to the nearest million

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

What's the exponential growth formula?

OpenStudy (anonymous):

P=Ae^kt

OpenStudy (whpalmer4):

Okay, if we take 1995 as t = 0, what value would we use for t for 2001 and 2010, respectively?

OpenStudy (anonymous):

2001=6? and 2010+15? Yes?

OpenStudy (whpalmer4):

Yes, I think so. What will we use for our value for \(A\)?

OpenStudy (anonymous):

2010=15*

OpenStudy (anonymous):

A is Number of the Population at Time Which The Time Is 0?

OpenStudy (whpalmer4):

right. and that is?

OpenStudy (anonymous):

228

OpenStudy (whpalmer4):

yes, if we're dealing in millions. That gives us \[P = 228 e^{kt}\]and we know that at \(t = 6, P = 230\) Can you solve for the value of \(k\)?

OpenStudy (anonymous):

K Is A Positive Constant..Thats What I Put In My Notes.So A Constant Would Be,I Honestly Dont Know?

OpenStudy (whpalmer4):

I tried punching "I Honestly Dont Know" into my calculator to find the population in 2010, but it just gave a funny look and beeped :-)

OpenStudy (whpalmer4):

Do you know about logarithms and exponentials?

OpenStudy (anonymous):

A Bit,Its Just That Im Doing A Online Course Of Algebra So I Dont Take It As Math Next Year..But Ive Never Learned These Type Of Equations..

OpenStudy (whpalmer4):

Okay, we'll do a quick demonstration. A logarithm is the inverse of an exponential. The logarithm (base 10) of \(x\) is the number \(n\) such that \(10^n = x\). That means the logarithm (base 10) of 100 would be 2, for example. Logs have some handy properties, such as: \[\log (a*b) = \log(a) + \log(b)\]\[\log(a/b) = \log(a)-\log(b)\]\[\log(a^b) = b\log(a)\] \[e \approx 2.71828 \text{ is the base of the natural logarithm, which is often denoted} \ln\] The natural logarithm is very handy in many scientific and mathematical contexts for reasons I won't try to cover, you'll just have to trust me :-) To solve this equation, all you need to know is how to punch the buttons on a scientific calculator or a website that offers similar functionality. Here's how: \[P = 228 e^{kt}\]\[230 = 228 e^{k*6}\]Divide both sides by 228\]\[\frac{230}{228} = e^{6k}\]Now we'll take the natural logarithm of both sides \[\ln \frac{230}{228} = 6k * \ln e\](by the last property I listed) and because the logarithm of the base of the logarithm is 1 (x^1 = x, right?) that gives us \[\ln \frac{230}{228} = 6k\]\[\frac{1}{6} \ln\frac{230}{228} = k\]That's just a simple bit of button-pushing on the calculator, and the result is approximately 0.00145561. Our formula becomes \[P = 228 e^{0.00145561t}\] and now we just plug in \(t=15\) to get our answer for the estimated population in 2010.

OpenStudy (whpalmer4):

If you have a scientific calculator, it probably has an e^x and ln x button pair. If not, you could use http://www.wolframalpha.com which will allow you to type it in like this: http://www.wolframalpha.com/input/?i=228+e%5E%280.00145561t%29%2C+t+%3D+15

OpenStudy (anonymous):

Woah. I Got 233..After I Round?correct?

OpenStudy (whpalmer4):

Yes, that's what I get as well. You should try the formula at t = 6 to check our work on finding \(k\). If we did it right, you should get 230.

OpenStudy (anonymous):

Yes Sir..Works Out Correct When I Plugged In and Checked..Thanks You So Much..and You Sir Have A Wonderful Day.. :)